how do i find out what is 2/5 of 5100 in percentage form
what is 30% of 5100

Answer by josmiceli(19441)

Always remember, percent means "per 100" , So you need to
know
2/5 = 40/100. This is 40%
2/5 of 5100 is the same as 40% of 5100
30% of 5100 = .3*5100
.3*5100 = 1530 answer